OpenStudy (anonymous):

MIkeeeeeeee12222 can somebody please help i dont understand thiss Suppose you have just enough money, in coins, to pay for a gallon of milk priced at $2.40. You have 12 coins, all quarters and dimes. Let q equal the number of quarters and d equal the number of dimes. Which system models the given information?

OpenStudy (solomonzelman):

Okay, so \(\normalsize\color{blue}{ \rm { q+d}=12 }\) this is the equation of number of coins \(\normalsize\color{blue}{ \rm { 0.25q+0.10d}=2.40 }\) and this is for the cost of the milk.

OpenStudy (solomonzelman):

Makes sense, right ?

OpenStudy (anonymous):

pretty much @SolomonZelman

OpenStudy (solomonzelman):

Okay;) Well.... Ill post something long. Ready ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{red}{\rm {Also,~~~for~~~a~~~future~~~reference:} }\) \(\normalsize\color{blue}{\rm {you~~~can~~~find~~~any~~~section~~~by:} }\) 1) clicking "more subjects" (the top left corner, towards the center) 2) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying). \(\normalsize\color{blue}{\rm {You~~~can~~~FAN~~~someone~~~by:} }\) 1) clicking on the user's name 2) clicking "fan" on the bottom right corner of user's box. ( you can also write a testimonial by clicking "add testimonial" and typing something in testimonial box) \(\normalsize\color{blue}{\rm {You~~~can~~~mention~~~someone~~~in~~~your~~~question} }\) (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @LOVEVAL (they would get a notification that a user mentioned them in a reply) \(\normalsize\color{blue}{\rm {You~~~can~~~ message~~~someone~~~by:} }\) 1) clicking on the someone's username 2) clicking "Send Message" 3) type your message press submit, and wait till the messages submits. If you have any questions you can message me or anyone else on this website. We will be happy to respond !

OpenStudy (anonymous):

@SolomonZelman ok now can we get back to the math problem?

OpenStudy (solomonzelman):

YEs, sure...

OpenStudy (solomonzelman):

At first, do you understand where I am getting the \(\normalsize\color{green}{ \bf {q+d=12 } }\) ?

OpenStudy (solomonzelman):

Then, \(\normalsize\color{blue}{ \rm {\color{red}{ 0.25}q+\color{red}{ 0.10}d=12 } }\) I am multiplying the number of quarters times 0.25, and the number of dimes times 0.10, THIS GIVES ME the cost.

OpenStudy (anonymous):

it gives you the cost because each quarter is equals 25cents and each dime is equal to 10 cents and the 12 is the total of coins put all together right?

OpenStudy (solomonzelman):

Oh yeah... can you solve the system, or need a little assistance for that part ? Or you don't need to solve the system ?

OpenStudy (anonymous):

i need a little assistance

OpenStudy (solomonzelman):

With solving the system of equations ?

OpenStudy (anonymous):

im sorry im late but yes pleasee

OpenStudy (coconutjj):

Subsitution: Isolate d from the second equation \[d=\frac{ 2.4-0.25q }{0.1}\] Subsitute d into the first equation, then solve: \[q+\frac{ 2.4-0.25q }{ 0.1 }=12\] Solve for d... Then plug the value of d back into either equation(preferably the first one), then solve for q